A firm faces the demand function p = 190 − 0.6q and the total cost function TC = 40 + 30q + 0.4q2 
(a) What output will maximize profit? 
(b) What output will maximize total revenue? 
(c) What will the output be if the firm makes a profit of 4,760?
1
Expert's answer
2020-08-27T13:37:20-0400

a)


"MC=p"


"MC=30+0.8q"

"30+0.8q=190-0.6q"

"q=114"

b)


"TR=pq"

"TR=190q-0.6q^2"


"TR^\/=190-1.2q"


"q=158"

c)


"TR-TC=160q-q^2-40"

"160q-q^2-40=4760"

"q=185"